THE MAGELLANIC CLOUDS NEWSLETTER

An electronic publication dedicated to the Magellanic Clouds, and astrophysical phenomena therein

No. 86 — 2 April 2007

http://www.astro.keele.ac.uk/MCnews

Editors: Jacco van Loon and Snežana Stanimirović

Editorial

Dear Colleagues,

It is our pleasure to present the 86^{th} issue of the Magellanic Clouds Newsletter.

This issue’s very interesting contributions include an independent distance measurement to the LMC based on δ Scuti stars (McNamara, Clementini & Marconi); the detection of a hot halo surrounding the LMC, possibly driving a galactic wind (Lehner & Howk); the identification of Herbig Ae/Be stars in the Magellanic Bridge young tidal feature (Nishiyama et al.); doubt casted on an historic association of the Magellanic Clouds and the Milky Way (Besla et al.); a parallax measurement of a microlens towards the SMC, favouring its location in the galactic halo (Dong et al.); and a Science paper on models that explain the circum-burst medium surrounding SN 1987A.

We are delighted to be hosting IAU Symposium 256, ”The Magellanic System: Stars, Gas, and Galaxies”, at Keele University (UK) from 28 July to 1 August 2008. A first announcement will be issued by the time of the next newsletter, but if you would already like to contact us the e-mail is: iaus256@astro.keele.ac.uk

The next issue will be distributed on the 1^{st} of June; the deadline for contributions is the 31^{st} of May.

Editorially Yours,

Jacco van Loon and Snežana Stanimirović

Refereed Journal Papers

A δ Scuti distance to the Large Magellanic Cloud

D. H. McNamara^{1}, G. Clementini^{2} and M. Marconi^{3}

1Department of Physics and Astronomy, Brigham Young University, N283 ESC, Provo, UT 84602, USA 2INAF - Osservatorio Astronomico di Bologna, Via Ranzani 1, I-40127 Bologna, Italy 3INAF - Osservatorio Astronomico di Napoli, Via Moiariello 16, I-80131 Napoli, Italy

We present results from a well studied δ Scuti star discovered in the Large Magellanic Cloud (LMC). The absolute magnitude of the variable was determined from the Period-Luminosity (P-L) relation for Galactic δ Scuti stars and from the theoretical modeling of the observed B,V,I light curves with nonlinear pulsation models. The two methods give distance moduli for the LMC of 18.46 ± 0.19 and 18.48 ± 0.15, respectively, for a consistent value of the stellar reddening of E(B − V) = 0.08 ± 0.02. We have also analyzed 24 δ Scuti candidates discovered in the OGLE II survey of the LMC, and 7 variables identified in the open cluster LW 55 and in the galaxy disk by Kaluzny et al. (2003, 2006). We find that the LMC δ Scuti stars define a P-L relation whose slope is very similar to that defined by the Galactic δ Scuti variables, and yield a distance modulus for the LMC of 18.50 ± 0.22 mag. We compare the results obtained from the δ Scuti variables with those derived from the LMC RR Lyrae stars and Cepheids. The corresponding distance moduli are: δ Scuti stars 18.48 ± 0.02 mag (standard deviation of the weighted average of the three above solutions); RR Lyrae stars 18.49 ± 0.06 mag; and Cepheids 18.53 ± 0.02 mag. We have assumed an average color excess of E(B − V) = 0.08 ± 0.02 mag and a metallicity correction of −0.19[Fe/H] for both δ Scuti stars and Cepheids. Within the observational uncertainties, the three groups of pulsating stars yield very similar distance moduli. These moduli are all consistent with the “long” astronomical distance scale for the Large Magellanic Cloud.

Accepted for publication in AJ Available from astro-ph/0702107

Highly ionised plasma in the Large Magellanic Cloud: Evidence for outflows and a possible galactic wind

N. Lehner^{1} and J.C. Howk^{1}

1University of Notre Dame, USA

Based on an analysis of the interstellar highly ionised species C iv, Si iv, N v, and O vi observed in the FUSE and HST/STIS E140M spectra of four hot stars in the Large Magellanic Cloud (LMC), we find evidence for a hot LMC halo fed by energetic outflows from the LMC disk and even possibly an LMC galactic wind. Signatures for such outflows are the intermediate and high-velocity components (v_{LSR} > 100 km/s) relative to the LMC disk observed in the highand low-ion absorption profiles. The stellar environments produce strong, narrow (T < 20, 000 K) components of C iv and Si iv associated with the LMC disk; in particular they are likely signatures of H ii regions and expanding shells. Broad components are observed in the profiles of C iv, Si iv, and O vi with their widths implying hot, collisionally ionised gas at temperatures of a few times 100,000 K. There is a striking similarity in the O vi/C iv ratios for the broad LMC and high-velocity components, suggesting much of the material at v_{LSR} > 100 km/s is associated with the LMC. The velocity of the high-velocity component is large enough to escape altogether the LMC, polluting the intergalactic space between the LMC and the Milky Way. The observed high-ion ratios of the broad LMC and highvelocity components are consistent with those produced in conductive interfaces; such models are also favored by the apparent kinematically coupling between the high and the weakly ionised species.

Accepted for publication in MNRAS Available from astro-ph/0702427

Three clusters of the SMC from ACS/WFC HST archive data: NGC 265, K 29 and NGC 290 and their field population

Emanuela Chiosi^{1} and Antonella Vallenari^{1}

1INAF, Padova Observatory, Vicolo Osservatorio 5, Padova, Italy

We determine the age, metallicity and initial mass function of three clusters, namely NGC 265, K 29, NGC 290, located in the main body of the Small Magellanic Cloud. In addition, we derive the history of star formation in the companion fields. We make use of ACS/WFC HST archive data. For the clusters, the age and metallicity are derived fitting the integrated luminosity function with single synthetic stellar population by means of the χ^{2} minimization. For the companion fields, the history of star formation is derived using the χ^{2} minimization together with the downhillsimplex method.For the clusters we find the following ages and metallicities: NGC 265 has log(Age)= 8.5 ± 0.3 yr and metallicity 0.004 ± 0.003 (or [Fe/H]=−0.62); K 29 has log(Age)=8.2 ± 0.2 yr and metallicity Z=0.003 ± 0.002 (or [Fe/H]=−0.75); NGC 290 has log(Age)=7.8 ± 0.5 yr and metallicity 0.003 ± 0.002(or [Fe/H]=−0.75). The superior quality of the data allows the study of the initial mass function down to M∼ 0.7 M_{⊙}. The initial mass function turns out to be in agreement with the standard Kroupa model. The comparison of the NGC 265 luminosity function with the theoretical ones from stellar models both taking overshoot from the convective core into account and neglecting it, seems to suggest that a certain amount of convective overshoot is required. However, this conclusion is not a strong one because this cluster has a certain amount of mass segregation which makes it difficult to choose a suitable area for this comparison. The star formation rate of the field population presents periods of enhancements at 300-400 Myr, 3-4 Gyr and finally 6 Gyr. However it is relatively quiescent at ages older than 6 Gyr. This result suggests that at older ages, the tidal interaction between the Magellanic Clouds and the Milky Way was not able to trigger significant star formation events.

Accepted for publication in A&A Available from astro-ph/0702281

Two SMC Symbiotic stars undergoing steady hydrogen burning

Orio, M.^{1}, Zezas, A.^{2}, Munari, U.^{3}, Siviero, A.^{4} and Tepedelenlioglu^{5}

1INAF-Padova, Italy, and Dept of Astronomy, University of Wisconsin, USA 2Center for Astrohysics, USA 3INAF-Padova (Osservatorio di Asiago), Italy 4Dept. of Astronomy, Padova University, Italy 5Physics Dept., University of Wisconsin, USA

Two symbiotic stars in the Small Magellanic Cloud (SMC), Lin 358 and SMC 3, have been supersoft X-ray sources (SSS) for more than 10 years. We fit atmospheric and nebular models to their X-ray, optical and UV spectra obtained at different epochs. The X-ray spectra are extremely soft, and appear to be emitted by the white dwarf atmosphere and not by the nebula like in some other symbiotics. The white dwarf of SMC 3, the hottest of the two sources, had a constant effective temperature ∼500,000 K at various epochs during 12 years. No nova-like outbursts of these systems have been recorded in the last 50 years, despite continuous optical monitoring of the SMC, and there are no indications of cooling of the white dwarf, expected after a thermonuclear flash. The bolometric luminosity of this system in March of 2003 was more than an order of magnitude lower than three years later, however the time of the observation is consistent with a partial eclipse of the white dwarf, previously found in ROSAT and optical observations. The red giant wind is either very asymmetric or very clumpy. The conpact object of Lin 358 has been at T > ∼180, 000 K since 1993, perhaps with a moderate increase. Atmospheric fits are obtained with log(g) ≥ 9, appropriate only for WD mass > 1.18 M_{⊙}. The two systems are probably accreting and burning hydrogen steadily at the high rate required for type Ia supernova progenitors.

Accepted for publication in ApJ Available from astro-ph/0702419

Herbig Ae/Be Stars in the Magellanic Bridge

Shogo Nishiyama^{1,2}, Yasuaki Haba^{2}, Daisuke Kato^{2}, Daisuke Baba^{2}, Hirofumi Hatano^{2}, Motohide Tamura^{1}, Yasushi Nakajima^{1}, Akika Ishihara^{1,3}, Tetsuya Nagata^{4}, Koji Sugitani^{5}, Noriyuki Matsunaga^{6}, Hinako Fukushi^{6}, Nobuhiko

Kusakabe^{7} and Shuji Sato^{2}

1National Astronomical Observatory of Japan, Japan 2Department of Astrophysics, Nagoya University, Japan 3Department of Earth and Planetary Science, School of Science, University of Tokyo, Japan 4Department of Astronomy, Kyoto University, Japan 5Graduate School of Natural Sciences, Nagoya City University, Japan 6Institute of Astronomy, School of Science, University of Tokyo, Japan 7Department of Astronomical Sciences, Graduate University for Advanced Studies (Sokendai), Japan

We have found Herbig Ae/Be star candidates in the western region of the Magellanic Bridge. Using the near infrared camera SIRIUS and the 1.4 m telescope IRSF, we surveyed about 3.0^{∘}×1.3^{∘} (24^{∘} < RA < 36^{∘}, −75^{∘} < Dec. < −73.7^{∘}) in the J, H, and K_{s} bands. On the basis of colors and magnitudes, about 200 Herbig Ae/Be star candidates are selected. Considering the contaminations by miscellaneous sources such as foreground stars and early-type dwarfs in the Magellanic Bridge, we estimate that about 80 (about 40%) of the candidates are likely to be Herbig Ae/Be stars. We also found one concentration of the candidates at the young star cluster NGC 796, strongly suggesting the existence of pre-main-sequence (PMS) stars in the Magellanic Bridge. This is the first detection of PMS star candidates in the Magellanic Bridge, and if they are genuine PMS stars, this could be direct evidence of recent star formation. However, the estimate of the number of Herbig Ae/Be stars depends on the fraction of classical Be stars, and thus a more precise determination of the Be star fraction or observations to differentiate between the Herbig Ae/Be stars and classical Be stars are required.

Accepted for publication in ApJ Available from astro-ph/0702504

First detection of phase-dependent colliding wind X-ray emission outside the Milky Way

Y. Naze^{1}, M.F. Corcoran^{2}, G. Königsberger^{3} and A.F.J. Moffat^{4}

1IAGL, University of Liège, Belgium 2GSFC, USA 3UNAM, México 4University of Montreal, Canada

After having reported the detection of X-rays emitted by the peculiar system HD5980, we assess here the origin of this high-energy emission from additional X-ray observations obtained with XMM-Newton. This research provides the first detection of apparently periodic X-ray emission from hot gas produced by the collision of winds in an evolved massive binary outside the Milky Way. It also provides the first X-ray monitoring of a Luminous Blue Variable only years after its eruption and shows that the dominant source of the X-rays is not associated with the ejecta.

Accepted for publication in ApJ (letters) Available from astro-ph/0702403

Are the Magellanic Clouds on their First Passage about the Milky Way?

Gurtina Besla^{1}, Nitya Kallivayalil^{1}, Lars Hernquist^{1}, Brant Robertson^{2,3,4}, T.J. Cox^{1}, Roeland P. van der Marel^{5} and Charles Alcock^{1}

1Harvard-Smithsonian CfA, USA 2KICP, U. Chicago, USA 3Enrico Fermi Institute, USA 4Spitzer Fellow 5STScI, USA

Recent proper motion measurements of the Large and Small Magellanic Clouds (LMC and SMC, respectively) by Kallivayalil (2006a,b) suggest that the 3D velocities of the Clouds are substantially higher (∼100 km/s) than previously estimated and now approach the escape velocity of the Milky Way (MW). Previous studies have also assumed that the Milky Way can be adequately modeled as an isothermal sphere to large distances. Here we re-examine the orbital history of the Clouds using the new velocities and a ΛCDM-motivated MW model with virial mass M_{vir} = 10^{12}M_{⊙} (e.g., Klypin et al. 2002). We conclude that the L/SMC are either currently on their first passage about the MW or, if the MW can be accurately modeled by an isothermal sphere to distances >200 kpc (i.e., M_{vir} > 2 × 10^{12}M_{⊙}), that their orbital period and apogalacticon distance must be a factor of two larger than previously estimated, increasing to 3 Gyr and 200 kpc, respectively. A first passage scenario is consistent with the fact that the LMC and SMC appear to be outliers when compared to other satellite galaxies of the MW: they are irregular in appearance and are moving faster. We discuss the implications of this orbital analysis for our understanding of the star formation history, the nature of the warp in the MW disk and the origin of the Magellanic Stream (MS), a band of HI i gas trailing the LMC and SMC that extends ∼100 degrees across the sky. Specifically, as a consequence of the new orbital history of the Clouds, the origin of the MS may not be explainable by current tidal and ram pressure stripping models.

Submitted to Astrophysical Journal Available from astro-ph/0703196

Efficiency of mass transfer in massive close binaries Tests from double-lined eclipsing binaries in the SMC

S. E. de Mink^{1}, O.R. Pols^{1} and R.W. Hilditch^{2}

1Astronomical Institute, Utrecht University, PO Box 80000, NL-3508 TA Utrecht, The Netherlands, S.E.deMink[AT]astro.uu.nl 2School of Physics and Astronomy. University of St Andrews, North Haugh, St Andrews, Fife KY16 9SS, Scotland, U.K.

One of the major uncertainties in close binary evolution is the efficiency of mass transfer β: the fraction of transferred mass that is accreted by a secondary star. We attempt to constrain the mass-transfer efficiency for short-period massive binaries undergoing case A mass transfer. We present a grid of about 20,000 detailed binary evolution tracks with primary masses 3.5–35 M_{⊙}, orbital periods 1–5 days at a metallicity Z = 0.004, assuming both conservative and non-conservative mass transfer. We perform a systematic comparison, using least-squares fitting, of the computed models with a sample of 50 double-lined eclipsing binaries in the Small Magellanic Cloud, for which fundamental stellar parameters have been determined. About 60% of the systems are currently undergoing slow mass transfer. In general we find good agreement between our models and the observed detached systems. However, for many of the semi-detached systems the observed temperature ratio is more extreme than our models predict. For the 17 semidetached systems that we are able to match, we find a large spread in the best fitting mass-transfer efficiency; no single value of β can explain all systems. We find a hint that initially wider systems tend to fit better to less conservative models. We show the need for more accurate temperature determinations and we find that determinations of surface abundances of nitrogen and carbon can potentially constrain the mass-transfer efficiency further.

Accepted for publication in A&A Available from astro-ph/0703480 and from http://www.astro.uu.nl/data/stars/

First Space-Based Microlens Parallax Measurement: Spitzer Observations of OGLE-2005-SMC-001

Subo Dong^{1,2}, A. Udalski^{3,4}, A. Gould^{1,2}, W.T. Reach^{5}, G.W. Christie^{2,6}, A.F. Boden^{7,8}, D.P. Bennett^{9}, G. Fazio^{10}, K. Griest^{11}, M.K. Szymański^{3,4}, M. Kubiak^{3,4}, I. Soszyński^{3,4}, G. Pietrzyński^{3,4,12}, O. Szewczyk^{3,4}, L. Wyrzykowski^{3,4,13}, K. Ulaczyk^{3,4}, T. Wieckowski^{3,4}, B. Paczyński^{3,14}, D. L. DePoy^{1,2}, R. W. Pogge^{1,2}, G. W.

Preston^{15}, I. B. Thompson^{15} and B. M. Patten^{10}

1Department of Astronomy, Ohio State University, 140 W. 18th Ave., Columbus, OH 43210, USA; depoy,dong,gould,pogge@astronomy.ohio-state.edu 2Microlensing Follow Up Network (μFUN) 3Optical Gravitational Lens Experiment (OGLE) 4Warsaw University Observatory, Al. Ujazdowskie 4, 00-478 Warszawa, Poland; udalski,msz,mk,soszynsk,pietrzyn,szewczyk,wyrzykow,kulaczyk,twieck@astrouw.edu.pl 5Spitzer Science Center, California Institute of Technology, Pasadena, CA 91125, USA; reach@ipac.caltech.edu 6Auckland Observatory, Auckland, New Zealand, gwchristie@christie.org.nz 7Michelson Science Center, California Institute of Technology, 770 South Wilson Ave., Pasadena, CA 91125, USA; bode@ipac.caltech.edu 8Department of Physics and Astronomy, Georgia State University, 29 Peachtree Center Avenue, Suite 400, Atlanta, GA 30303, USA. 9Department of Physics, Notre Dame University, Notre Dame, IN 46556, USA; bennett@nd.edu 10Center for Astrophysics, Cambridge, MA 02138, USA; fazio,bpatten@cfa.harvard.edu 11Department of Physics, University of California, San Diego, CA 92093, USA; griest@ucsd.edu 12Universidad de Concepción, Departamento de Fisica, Casilla 160–C, Concepción, Chile 13Institute of Astronomy Cambridge University, Madingley Rd., CB3 0HA Cambridge, UK 14Princeton University Observatory, Princeton, NJ 08544, USA; bp@astro.princeton.edu 15The Observatories of the Carnegie Institute of Washington, 813 Santa Barbara Street, Pasadena, CA 91101, USA; gwp,ian@ociw.edu

We combine Spitzer and ground-based observations to measure the microlens parallax of OGLE-2005-SMC-001, the first space-based such determination since Refsdal proposed the idea in 1966. The parallax measurement yields a projected velocity ṽ ∼ 230 km s^{−1}, the typical value expected for halo lenses, but an order of magnitude smaller than would be expected for lenses lying in the Small Magellanic Cloud itself. The lens is a weak (i.e., non-caustic-crossing) binary, which complicates the analysis considerably but ultimately contributes additional constraints. Using a test proposed by Assef et al. (2006), which makes use only of kinematic information about different populations but does not make any assumptions about their respective mass functions, we find that the likelihood ratio is L_{halo}/L_{SMC} = 20. Hence, halo lenses are strongly favored but SMC lenses are not definitively ruled out. Similar Spitzer observations of additional lenses toward the Magellanic Clouds would clarify the nature of the lens population. The Space Interferometry Mission could make even more constraining measurements.

Accepted for publication in The Astrophysical Journal Available from astro-ph/0702240

The Triple-Ring Nebula Around SN 1987A: Fingerprint of a Binary Merger

Thomas Morris^{1,2} and Philipp Podsiadlowski^{1}

1Department of Astrophysics, University of Oxford, Oxford OX1 3RH, UK 2Max-Planck Institut für Astrophysik, Garching 85741, Germany

Supernova 1987A, the first naked-eye supernova observed since Kepler’s supernova in 1604, defies a number of theoretical expectations. Its anomalies have long been attributed to a merger between two massive stars that occurred some 20,000 years before the explosion, but so far there has been no conclusive proof that this merger took place. Here, we present three-dimensional hydrodynamical simulations of the mass ejection associated with such a merger and the subsequent evolution of the ejecta, and we show that this accurately reproduces the properties of the triple-ring nebula surrounding the supernova.

Published in Science Available from astro-ph/0703317

Conference Papers

The VMC survey and the SFH of some Local Group Galaxies

Maria-Rosa L. Cioni^{1}

1SUPA, School of Physics, University of Edinburgh, Edinburgh, EH9 3HJ, UK

The K_{s}-band magnitude distribution of carbon-rich and oxygen-rich asymptotic giant branch stars within Local Group galaxies like the Magellanic Clouds, NGC 6822, M33 and SagDIG is easily obtained from ground-based observations. Appropriate stellar evolutionary models covering a range of metallicities and star formation rates are used to produce theoretical distributions that allow us to derive the history of star formation across these galaxies. I will show the result of these studies and discuss the application of this technique to more distant systems as well as deeper observations, like those that VISTA will provide, to improve our understanding of, in particular, the Magellanic Clouds.

Oral contribution, published in IAU Symp. 241, Stellar Populations as Building Blocks of Galaxies Available from astro-ph/0701910

Old main-sequence turnoff photometry in the SMC: Star Formation History and Chemical Enrichment Law

Noelia E. D. Noël^{1}, Carme Gallart^{1}, Antonio Aparicio^{1}, Sebastián L. Hidalgo^{2}, Edgardo Costa^{3} and René A. Méndez^{3}

1Instituto de Astrofísica de Canarias, Spain 2University of Minnesota, USA 3Universidad de Chile, Chile

We present deep ground-based B and R observations of 12 fields in the Small Magellanic Cloud (SMC). The resulting color-magnitude diagrams (CMDs) reach the oldest main-sequence (MS) turnoff at M_{R} ∼3.5 and reveal the stellar population differences between the part of the galaxy facing the Large Magellanic Cloud (LMC) and an area on the opposite side. In the Southern part of the galaxy, we found that there are still intermediate-age stars as far as 4 kpc from the SMC center. The Chemical Enrichment History (CEH) in one of our SMC fields is also presented.

Oral contribution, published in Stellar Populations as Building Blocks of Galaxies, Proceedings IAU Symposium No. 241, 2007 Available from astro-ph/0703225

The Star Formation History in a SMC field: IAC-star/IAC-pop at work

Noelia E. D. Noël^{1}, Antonio Aparicio^{1}, Carme Gallart^{1}, Sebastián L. Hidalgo^{2}, Edgardo Costa^{3} and René A. Méndez^{3}

1Instituto de Astrofísica de Canarias, Spain 2University of Minnesota, USA 3Universidad de Chile, Chile

We present a progress report of a project to study the quantitative star formation history (SFH) in different parts of the Small Magellanic Cloud (SMC). We use the information in [(B-R), R] color-magnitude diagrams (CMDs), which reach down to the oldest main-sequence turnoffs and allow us to retrieve the SFH in detail. We show the first results of the SFH in a SMC field located in the Southern direction (at ∼1 kpc from the SMC center). This field is particularly interesting because in spite of being located in a place in which the H i column density is very low, it still presents a recent enhancement of star formation.

Poster contribution, published in Stellar Populations as Building Blocks of Galaxies, Proceedings IAU Symposium No. 241, 2007 Available from astro-ph/0703222

Limits on iron-dominated fallback disk in SN 1987A

K. Werner^{1}, T. Nagel^{1} and T. Rauch^{1}

1Institut für Astronomie und Astrophysik, University of Tübingen, Germany

The non-detection of a point source in SN1987A imposes an upper limit for the optical luminosity of L = 2 L_{⊙}. This limits the size of a possible fallback disk around the stellar remnant. Assuming a steady-state thin disk with blackbody emission requires a disk smaller than 100,000 km if the accretion rate is at 30% of the Eddington rate (Graves et al. 2005). We have performed detailed non-LTE radiation transfer calculations to model the disk spectrum more realistically. It turns out that the observational limit on the disk extension becomes even tighter, namely 70,000 km.

Oral contribution, published in Supernova 1987A: Twenty years after Available from astro-ph/0703520 and from http://astro.uni-tuebingen.de/publications/paper 07 05.shtml

ALBUM: a tool for the analysis of slitless spectra and its application to ESO WFI data.

C. Martayan^{1,2}, D. Baade^{1}, A.-M. Hubert^{2}, M. Floquet^{2}, J. Fabregat^{3} and E. Bertin^{4}

1European Organisation for Astronomical Research in the Southern Hemisphere, Karl-Schwarzschild-Str. 2, 85748 Garching b. München, Germany 2GEPI-Observatoire de Paris, 5 place Jules Janssen, 92195 Meudon cedex, France 3Observatori Astronòmic de la Universitat de València, Edifici Instituts d’Investigació, Polígon La Coma, 46980 Paterna València, Spain 4Institut d’Astrophysique de Paris, 98bis boulevard Arago, 75014 Paris, France

ALBUM is a general-purpose tool to visualize and screen large amounts of slitless spectra. It was developed for a search for emission-line stars in SMC and LMC clusters. The observations were obtained with ESO’s Wide Field Imager (WFI) and comprise ∼8 million low-resolution spectra. The tool as well as the results of its application to the SMC part of the database are presented. The inferred frequency of Be stars is compared to the one in the higher-metallicity environment of the Milky Way.

Poster contribution, published in ESO instrument calibration workshop 2007, to appear in the Springer-Verlag series Available from astro-ph/0703605